Required information Use the following information for the Exercises below. [The following information applies to the questions displayed below.] Hart Company made 4,500 bookshelves using 21,000 board feet of wood costing $241,500. The companys direct materials standards for one bookshelf are 5 board feet of wood at $11.40 per board foot. Exercise 8.13 C。mputation and interpretation of materials variances LO P2 1) Compute the direct materials price and quantity variances incurred in manufacturing these bookshelves. AQ- Actual Quantity SQ Standard Quantity AP= Actual Price SP Standard Price Actual Cost Standard Cost
Answer #1

Actual quantity (AQ)= 21000

Standard quantity (SQ)= 22500 (5*4500)

Actual price (AP)= 11.5 (241500/21000)

Standard price (SP) = 11.4

Direct material price variance = AQ*(SP-AP)

= 21000(11.4-11.5)

=21000* -0.1

= -2100 ( unfavorable)

Direct material quantity variance = SP*(SQ-AQ)

=11.4(22500-21000)

=11.4(1500)

= 17100 ( favourable)